# Quantising Chiral Bosons On Riemann Surfaces

### v2

@@ 2.1 The Action For Chiral Gauge Fields
Its role is to ensure that
@@
[…]
The equations of motion of this action take the form
@@ 2.2 The action For Chiral Scalars
and $G_{-}=0$, $F=\ast F$, $G_{+}=\partial_{+}C$, $F_{\pm}=\partial_{\pm}A$.
As discussed in the introduction, for a given metric $g$ we will choose $\bar{g}$ so that $M_{--}(x)$ is finite and non-zero for all $x$. We then find that the action (14) can be rewritten as
@@
$\hskip 96.73918pt-M_{--}(Q_{+}+J_{+}+M_{11}^{-1}(\partial_{-}P-J_{-}))^{2}$
$\hskip 96.73918pt+2(J_{+}-M_{11}^{-1}J_{-})\partial_{-}P-J_{+}J_{-}+M_{--}^{-1}J_{-}^{2}\Big{)}d^{2}x,$ (19)
with the notation $\partial_{\pm}P=\bar{e}_{\pm}{}^{\mu}\partial_{\mu}P$. Shifting $Q$ and $J$ to define
@@
Thus we see that the action splits into an effective action involving the scalar $P$ alone, coupled to sources, together with a simple quadratic action in the field $Q^{\prime}_{+}$. The action for $Q^{\prime}_{+}$ involves no derivatives, so that $Q^{\prime}_{+}$ is an auxiliary field.
@@ 2.4 Evaluating $S_{eff}$
We now discuss further the effective action (23) that we found for $P$. We write it as
@@ 3.2 A Canonical Approach
For this to be non-zero we must also introduce a periodicity on $P$; $P\cong P+2\pi R$. In this case we must have $a_{0}-\tilde{a}_{0}=2mR$ for some constant $R$ and integer $m\in\mathbb{Z}$. To this end, we write
$a_{0}=p_{0}+mR,\qquad\tilde{a}_{0}=p_{0}-mR.$ (76)
@@ 4.2 Gravitational Anomalies
As mentioned above, in the bi-metric chiral boson action, there are two gauge symmetries, one for which $g$ is the gauge field and one for which $\bar{g}$ is the gauge field, with the ‘true’ diffeomorphisms arising as a diagonal subgroup of these Hull:2023dgp . Anomalies arise for both these symmetries and these will be discussed more fully elsewhere.
@@ 5.3 Holomorphic Factorisation And Even Self-Dual Lattices
For the torus CFT’s, the sum over oscillator modes gives
@@
which factorises into a function of $\tau$ and a function of $\tau^{*}$, but the inclusion of momenta and winding from the zero modes in general gives a function of both $\tau$ and $\tau^{*}$. In the cases in which this a rational CFT, the partition function is a finite sum of terms, each of which factorises into a function of $\tau$ and a function of $\tau^{*}$.
For the Sen-type chiral boson action, the sum over oscillators gives a term which factorises into the product of a function of $\tau$ and a function of $\rho$, but now the momentum and winding contributions give a function of both $\tau$ and $\rho$ in general. In the rational case, this can be written as a finite sum of terms, each of which factorises into a function of $\tau$ times a function of $\rho$.
For certain lattices, there can be points in the Narain moduli space for which there is a complete factorisation
@@
for some functions $F,G$ and such points will have an important role here. In this case, the chiral partition function for chiral bosons can be taken to be
$Z(\tau)\sim F(\tau).$ (168)
Then for the corresponding Sen-type chiral boson theory the partition function is
$Z(\tau,\rho)\sim F(\tau)G(\rho).$ (169)
Consider a theory with $n$ chiral bosons with partition function $Z(\tau)\sim F(\tau)$ coupled to other fields which couple to $g$ but not $\bar{g}$ that have a partition function $f(\tau,\tau^{*})$ such that the total partition function
$Z(\tau,\tau^{*})\sim f(\tau,\tau^{*})F(\tau),$ (170)
is modular invariant. If the $n$ chiral bosons are replaced by the Sen-type chiral boson theory with partition function $Z(\tau,\rho)\sim F(\tau)G(\rho)$, then the partition function for the whole system is of the form
$Z(\tau,\rho)\sim f(\tau,\tau^{*})F(\tau)G(\rho),$ (171)
which is invariant under $\tau$ modular transformations. On the other hand, in the general case in which $Z(\tau,\rho)$ is a sum of terms that each factorise, then modular invariance would be hard to achieve.
Note that the partition function (171) will not in general be invariant under $\rho$ modular transformations. If $\bar{g}$ is treated as a fixed background field, then this may not be important. However, if we define a string-like theory in which we integrate over both $\bar{g}$ as well as the physical metric $g$, then $\rho$ modular invariance would be desirable. This could be achieved by adding further matter coupling to $\bar{g}$ but not $g$ with partition function $\tilde{f}(\rho,\rho^{*})$ such that $\tilde{f}(\rho,\rho^{*})G(\rho)$ is modular invariant. Then the total partition function
$Z(\tau,\rho)\sim f(\tau,\tau^{*})\tilde{f}(\rho,\rho^{*})F(\tau)G(\rho),$ (172)
would be invariant under separate $\tau$ and $\rho$ modular transformations.
Such a complete factorisation arises at rational points in the Narain moduli space at which $\Lambda=\Gamma_{L}\oplus\Gamma_{R}$ with $\Gamma_{L}=\Gamma_{R}=\Delta$ where $\Delta$ is an even self-dual lattice. This requires $n$ to be a multiple of $8$. The partition function is then of the form
$Z(\tau,\tau^{*})\sim F(\tau)F(\tau^{*}).$ (173)
@@ 5.4 The Heterotic String
The heterotic string can be formulated with 16 chiral bosons on $\mathbb{R}^{16}/\Delta$ where $\Delta$ is a 16-dimensional even self-dual lattice. There are precisely two such lattices: the root lattice of $E_{8}\times E_{8}$ and the weight lattice of $Spin(32)/\mathbb{Z}_{2}$. These have partition function

### v3

@@ 2.1 The Action For Chiral Gauge Fields
It is also taken to be symmetric in the sense that
$Q_{1}\wedge{\cal M}(Q_{2})=Q_{2}\wedge{\cal M}(Q_{1}),$ (9)
[…]
The role of ${\cal M}(Q)$ is to ensure that
@@
where $\star$ is the Hodge dual associated to a second, physical metric $g$ that is independent of $\bar{g}$. The equations of motion of the action (2.1) take the form
@@ 2.2 The action For Chiral Scalars
and $G_{-}=0$, $F=\ast F$.
As discussed in the introduction, for a given metric $g$ we will restrict our attention to choices of $\bar{g}$ so that $M_{--}(x)$ is well-defined and non-zero for all $x$. We then find that the action (15) can be rewritten as
@@
$\hskip 96.73918pt-M_{--}(Q_{+}+J_{+}+M_{--}^{-1}(\partial_{-}P-J_{-}))^{2}$
$\hskip 96.73918pt+2(J_{+}-M_{--}^{-1}J_{-})\partial_{-}P-J_{+}J_{-}+M_{--}^{-1}J_{-}^{2}\Big)d^{2}x,$ (20)
with the notation $\partial_{\pm}P=\bar{e}_{\pm}{}^{\mu}\partial_{\mu}P$. To clean up this action let us shift $Q$ and $J$ are define
@@
Thus we see that the action splits into two decoupled parts parts: a contribution involving the scalar $P$ alone, coupled to sources, together with a simple quadratic term in the field $Q^{\prime}_{+}$. In particular the action for $Q^{\prime}_{+}$ involves no derivatives, so that $Q^{\prime}_{+}$ is an auxiliary field and so can be integrated out to give a result that is independent of $P$ and the sources.
@@ 2.4 Evaluating $S_{eff}$
We now discuss further the effective action (24) that we found for $P$. It is helpful to write it in a more familiar form as
@@ 4.2 Gravitational Anomalies
As mentioned above, in the bi-metric chiral boson action, there are two gauge symmetries, one for which $g$ is the gauge field and one for which $\bar{g}$ is the gauge field, with the ‘true’ diffeomorphisms arising as a diagonal subgroup of these Hull:2023dgp . Anomalies arise for both these symmetries and these will be discussed more fully elsewhere (see Hull:2025bqo and future work).
@@ 5.3 Holomorphic Factorisation And Even Self-Dual Lattices
For a standard bosonic CFT the sum over oscillator modes gives a contribution to the partition function of the form
@@
which factorises into a function of $\tau$ times a function of $\tau^{*}$, but the inclusion of momenta and winding from the zero modes in general gives a function of both $\tau$ and $\tau^{*}$ that doesn’t factorise. For the special tori for which the CFT is rational, the partition function is a finite sum of terms, each of which factorises into a function of $\tau$ and a function of $\tau^{*}$.
For the Sen-type chiral boson action, the sum over oscillators gives a term which factorises into the product of a function of $\tau$ and a function of $\rho$, but now the momentum and winding contributions give a function of both $\tau$ and $\rho$ in general. Just as before, in the rational case, this can be written as a finite sum of terms, each of which factorises into a function of $\tau$ times a function of $\rho$.
For certain lattices, there can be points in the Narain moduli space for which there is a complete factorisation of the full partition function (including zero modes)
@@
for some functions $F,G$ and such points will have an important role here. Such a complete factorisation arises at rational points in the Narain moduli space at which $\Lambda=\Gamma_{L}\oplus\Gamma_{R}$ with $\Gamma_{L}=\Gamma_{R}=\Delta$ where $\Delta$ is an even self-dual lattice. This can only occur if $n$ is a multiple of $8$. The partition function is then of the form
$Z(\tau,\tau^{*})\sim F(\tau)F^{*}(\tau^{*}).$ (168)
@@
In these cases, the chiral partition function for chiral bosons can be taken to be
$Z(\tau)\sim F(\tau).$ (172)
Similarly for the corresponding Sen-type chiral boson theory the partition function is
$Z(\tau,\rho)\sim F(\tau)F^{*}(\rho).$ (173)
Next we consider a theory with $n$ chiral bosons with partition function $Z(\tau)\sim F(\tau)$ and embed it in a larger theory with other fields which couple to $g$ but not $\bar{g}$. For example we will consider the heterotic string in the next section. If the additional fields have a modular partition function $f(\tau,\tau^{*})$ then the total partition function
$Z(\tau,\tau^{*})\sim f(\tau,\tau^{*})F(\tau),$ (174)
is modular invariant. If the $n$ chiral bosons are replaced by the Sen-type chiral boson theory with partition function $Z(\tau,\rho)\sim F(\tau)F^{*}(\rho)$, then the partition function for the whole system is of the form
$Z(\tau,\rho)\sim f(\tau,\tau^{*})F(\tau)F^{*}(\rho),$ (175)
which is still invariant under $\tau$ modular transformations. On the other hand, in the general case in which $Z(\tau,\rho)$ is a sum of terms that each factorise, then modular invariance would be hard, if not impossible, to achieve.
Note that the partition function (175) may now have gravitational anomalies associated to $\bar{g}$. If $\bar{g}$ is treated as a fixed background field, then this may not be important. On the other hand, if we define a string-like theory in which we integrate over both $\bar{g}$ as well as the physical metric $g$, then $\rho$ modular invariance and no gravitational anomalies would be desirable. This could be achieved by adding further matter coupling to $\bar{g}$ but not $g$ with partition function $\tilde{f}(\rho,\rho^{*})$ such that $\tilde{f}(\rho,\rho^{*})F^{*}(\rho)$ is modular invariant and anomaly free. Then the total partition function
$Z(\tau,\rho)\sim f(\tau,\tau^{*})\tilde{f}(\rho,\rho^{*})F(\tau)F^{*}(\rho),$ (176)
would be invariant under separate $\tau$ and $\rho$ modular transformations. Let us now consider the more specific example of the heterotic string.
@@ 5.4 The Heterotic String
The heterotic string can be formulated with 16 chiral bosons on $\mathbb{R}^{16}/\Delta$ where $\Delta$ is a 16-dimensional even self-dual lattice. There are precisely two such lattices: the root lattice of $E_{8}\times E_{8}$ and the weight lattice of $Spin(32)/\mathbb{Z}_{2}$. These both have the partition function